OpenStudy (highschoolmom2010):

determine whether you would use the Law of Sines or Law of Cosines to find angle x. Then find angle x to the nearest tenth. 8.8 by the law of sine (SSA) 22.9 by the law of cosine (SAS) 10.7 law of sine (SAS) 68.2 by law of cosine (SSS)

OpenStudy (highschoolmom2010):



OpenStudy (john_es):

In this problem you need length X, no angle X. You should use the law of sines, because you have two angles and one side opposite to one of the given angles. So, \[\frac{X}{\sin 21}=\frac{27}{\sin 65}\Rightarrow X=10.7\]

OpenStudy (highschoolmom2010):

so now I do x sin (65)=27 si(n 21)=10.7 thanks :D @John_ES

OpenStudy (john_es):

Exact ;).